OpenStudy (anonymous):

please help will give medal!!!! When given all three side lengths of a triangle but none of the angle measures, you can solve for all angle measures using __________.

OpenStudy (schrodingers_cat):

Well if you use the law of cosines you can solve for an angle if you know all three sides. Like this c2 = a2 + b2 - 2ab cos(C) Cos(c) = (a2+b2 -c2)/2ab Hope this helps :)